OpenStudy (anonymous):

if the measure of an exterior angle of a regular polygon is 40°, how many sides does the polygon have?

OpenStudy (ranga):

All he external degrees of a polygon add up to 360 degrees. If there are N sides, there will be N external angles. N * 40 = 360 N = ?

OpenStudy (anonymous):

An equilateral triangle has 3 interior angles of 60., exterior f 120o each totallng 360o exterior. Square has four 90o exterior angles, total = 360. Looks like total of exterior angles is 360 so angle = 360/ n sides. n=9?

OpenStudy (anonymous):

Thanks, @ranga. I was just speculating.